On July 1, 2021, Ross-Livermore Industries issued nine-month notes in the amount of $600 million. Interest is payable at maturity.

Required:

Determine the amount of interest expense that should be recorded in a year-end adjusting entry under each of the following independent assumptions: (Enter your answers in millions (i.e., 10,000,000 should be entered as 10).)

Fiscal Year End: Principal (million) x Interest Rate x Time = Interest Expense
December 31, 2021 $600 x 10 % x = million
September 30, 2021 $600 x 8 % x = million
October 31, 2021 $600 x 7 % x = million
January 31, 2022 $600 x 6 % x = million
Answer #1

Answer

Fiscal Year End: Principal (million) x Interest Rate x Time = No of months since 1st July to Year end = Interest Expense
December 31, 2021 $600 x 10 % x 6 / 12 = $30 million
September 30, 2021 $600 x 8 % x 3 / 12 = $12 million
October 31, 2021 $600 x 7 % x 4 / 12 = $14 million
January 31, 2022 $600 x 6 % x 7 / 12 = $21 million